NatalieAfton:

Finish solving the system of equations, y = x – 15 and y = –2x + 3, using the substitution method. 1. Use substitution to create a one-variable linear equation:   x – 15 = –2x + 3 2. Solve to determine the unknown variable in the equation: 3x = 18 x = 6 3. Substitute the value of the variable into either original equation to solve for the other variable. 4. Write the solution to the system of equations as an ordered pair. The solution to the system is-? (9,6) (9,-6) (6,9) (6,-9)

jimthompson5910:

You'll use x = 6 in either y = x-15 or y = -2x+3 to find y.

jimthompson5910:

As a similar example, let's say we know that x = 5 and we want to use it to find y when y = 2x+7 That would mean, y = 2x+7 y = 2*5+7 ... replace x with 5 y = 10+7 y = 17 The solution to this example would be (x,y) = (5,17)

NatalieAfton:

?

NatalieAfton:

T^T

jimthompson5910:

Which part are you stuck on?

NatalieAfton:

(6,-9)

jimthompson5910:

Yes that's correct because y = x-15 y = 6-15 ... x replaced with 6 y = -9 or we could say y = -2x+3 y = -2(6) + 3 .... x replaced with 6 y = -12+3 y = -9 Either way, we get the same y value. This helps us confirm that x = 6 is correct.

NatalieAfton:

k

NatalieAfton:

thx

ZombieLover:

y = -2x+3 y=-2*6+3 (6*2) y=12 +3 12+3= ? (equals x)

yourgrandpa:

if someone answered already you can close this question to ask another one (:

NatalieAfton:

thx for giving me the answer brooooooo!!

NatalieAfton:

yourgrandpa is my bro!

ZombieLover:

lol

NatalieAfton:

lol

yourgrandpa:

lol sis